OpenStudy (anonymous):

how to find the ordered pairs of x=7-y

jimthompson5910 (jim_thompson5910):

pick any value you want let's say you pick 0 if you replace y with 0 you get x = 7 - y x = 7 - 0 x = 7 which means that the ordered pair (7,0) is on this line

jimthompson5910 (jim_thompson5910):

pick a few other values to plug into y to get other ordered pairs

OpenStudy (anonymous):

x=7-3

jimthompson5910 (jim_thompson5910):

so if y = 3, then x = 4 (since 7-3 = 4)

jimthompson5910 (jim_thompson5910):

another ordered pair is (4,3)